The sum of two integers is 15 and their product is 54.What is the quotient, in
simplest form, when the smaller integer is divided by the larger one?

Answer by mananth(16949)

x+y=15
xy=54
..
x=54/y
54/y +y=15
54+y^2 =15y
y^2-15y+54=0
y^2-9x-6x+54=0
y(y-9)-6(y-9)=0
(y-6)(y-9)=0
y=6 or 9
the numbers are 6 & 9
6/9 = 2/3